# Probing Primordial Black Hole Formation from Domain Wall Isocurvature Perturbations: Constraints and Implications

### v2

@@ Appendix E Strategy for PTA
In this section, we provide further details of the PTA data analysis. We follow the analysis procedure encoded in PTArcade, which wraps enterprise and Ceffyl to allow us to perform a standard search strategy for the PTA datasets (see Ref. [93] for more details). The searches for PTA stochastic signals adopt the pulsars’ timing residuals, which are modeled as
@@
where the terms on the right-hand side of Eq. (E.1) are the white noise, red noise, and small errors in the fit to the timing-ephemeris parameter [115], respectively.
@@
$E_{k}$ EFAC per backend/receiver system Uniform [0, $0,\leavevmode\nobreak\ 1$0] Single-pulsar analysis only
$Q_{k}$[s] EQUAD per backend/receiver system Log-uniform $8.5,\leavevmode\nobreak\ -5]$ Single-pulsar analysis only
$J_{k}$[s] ECORR per backend/receiver system Log-uniform $8.5,\leavevmode\nobreak\ -5]$ Single-pulsar analysis only
@@
$A_{\rm red}$ Red noise power-law amplitude Log-uniform $20,\leavevmode\nobreak\ -11]$ one parameter per pulsar
$\gamma_{\rm red}$ Red noise power-law spectral index Uniform $0,\leavevmode\nobreak\ 7]$ one parameter per pulsar
@@
$T_{\rm ann}$ Domain wall annihilation temperature Log-Uniform $2.3,\leavevmode\nobreak\ 0]$ one parameter per PTA
$c_{1}$ GW spectral index above peak frequency Uniform $0.5,\leavevmode\nobreak\ 1.5]$ one parameter per PTA
$c_{2}$ Width of spectral index Uniform $0.5,\leavevmode\nobreak\ 2.5]$ one parameter per PTA
@@ Appendix F Models for $N$ domain walls
To realize a model with $\mathcal{N}$ domains (and the domain wall number $N\sim\mathcal{N}$), we can consider the model with a $U(1)$ symmetry, which is softly broken by the term [21, 116]
@@
One can consider $\mathcal{N}$ copies of complex scalars $\Phi_{i}$ (with $i=1,...,\mathcal{N}$). The global $U(1)^{\mathcal{N}}$ symmetry is spontaneously broken when each of $\Phi_{i}$ obtains a common vacuum expectation value. We can further explicitly break the $\mathcal{N}$ $U(1)$ symmetries to $\mathcal{N}-1$ discrete shift symmetries and one $U(1)$ symmetry by a term $\Phi_{i}^{\dagger}\Phi_{i+1}^{q}$, which then leads to the formation of $\mathcal{N}-1$ domains. Such symmetries have been naturally embedded in the framework of the clockwork axion models [117] and deconstruction of extradimensional grand unified theories (GUTs) [118, 119].
Another realization of the $Z_{\mathcal{N}}$ symmetry is the axion-like models (see, for example, Ref. [120]). Consider a hidden non-Abelian $SU(n)$ gauge sector consisting of a complex scalar $\Phi$ and a set of vector-like quarks. The potential of $\Phi$ possesses a global $U(1)$ symmetry, which is spontaneously broken by the Higgs mechanism. If the vector-like quarks are charged under the $U(1)$ transformation, then the quantum anomaly of $U(1)$ with respect to the $SU(n)$ gauge group can break the $U(1)$ symmetry to its $Z_{\mathcal{N}}$ subgroup via the instanton effect during the confinement of the $SU(n)$ symmetry, which gives rise to an axion-like periodic potential

### v3

@@ Appendix E Strategy for PTA
In this section, we provide further details of the PTA data analysis. We follow the analysis procedure encoded in PTArcade, which wraps enterprise and Ceffyl to allow us to perform a standard search strategy for the PTA datasets (see Ref. [Mitridate:2023oar] for more details). The searches for PTA stochastic signals adopt the pulsars’ timing residuals, which are modeled as
@@
where the terms on the right-hand side of Eq. (E.1) are the white noise, red noise, and small errors in the fit to the timing-ephemeris parameter [NANOGrav:2020tig], respectively.
@@
$E_{k}$ EFAC per backend/receiver system Uniform [0, $0,\penalty 10000\ 1$0] Single-pulsar analysis only
$Q_{k}$[s] EQUAD per backend/receiver system Log-uniform $8.5,\penalty 10000\ -5]$ Single-pulsar analysis only
$J_{k}$[s] ECORR per backend/receiver system Log-uniform $8.5,\penalty 10000\ -5]$ Single-pulsar analysis only
@@
$A_{\rm red}$ Red noise power-law amplitude Log-uniform $20,\penalty 10000\ -11]$ one parameter per pulsar
$\gamma_{\rm red}$ Red noise power-law spectral index Uniform $0,\penalty 10000\ 7]$ one parameter per pulsar
@@
$T_{\rm ann}$ Domain wall annihilation temperature Log-Uniform $2.3,\penalty 10000\ 0]$ one parameter per PTA
$c_{1}$ GW spectral index above peak frequency Uniform $0.5,\penalty 10000\ 1.5]$ one parameter per PTA
$c_{2}$ Width of spectral index Uniform $0.5,\penalty 10000\ 2.5]$ one parameter per PTA
@@ Appendix F Models for $N$ domain walls
To realize a model with $\mathcal{N}$ domains (and the domain wall number $N\sim\mathcal{N}$), we can consider the model with a $U(1)$ symmetry, which is softly broken by the term [Chiang:2019oms, Bai:2023cqj]
@@
One can consider $\mathcal{N}$ copies of complex scalars $\Phi_{i}$ (with $i=1,...,\mathcal{N}$). The global $U(1)^{\mathcal{N}}$ symmetry is spontaneously broken when each of $\Phi_{i}$ obtains a common vacuum expectation value. We can further explicitly break the $\mathcal{N}$ $U(1)$ symmetries to $\mathcal{N}-1$ discrete shift symmetries and one $U(1)$ symmetry by a term $\Phi_{i}^{\dagger}\Phi_{i+1}^{q}$, which then leads to the formation of $\mathcal{N}-1$ domains. Such symmetries have been naturally embedded in the framework of the clockwork axion models [Kaplan:2015fuy] and deconstruction of extradimensional grand unified theories (GUTs) [Arkani-Hamed:2001kyx, Li:2002xd].
Another realization of the $Z_{\mathcal{N}}$ symmetry is the axion-like models (see, for example, Ref. [ZambujalFerreira:2021cte]). Consider a hidden non-Abelian $SU(n)$ gauge sector consisting of a complex scalar $\Phi$ and a set of vector-like quarks. The potential of $\Phi$ possesses a global $U(1)$ symmetry, which is spontaneously broken by the Higgs mechanism. If the vector-like quarks are charged under the $U(1)$ transformation, then the quantum anomaly of $U(1)$ with respect to the $SU(n)$ gauge group can break the $U(1)$ symmetry to its $Z_{\mathcal{N}}$ subgroup via the instanton effect during the confinement of the $SU(n)$ symmetry, which gives rise to an axion-like periodic potential